OpenStudy (anonymous):

David's company reimburses his expenses on food, lodging, and conveyance during business trips. The company pays $60 a day for food and lodging and $0.65 for each mile traveled. David drove 600 miles and was reimbursed $3390. Part A: Create an equation that will determine the number of days x on the trip (3 points) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) Part C: How many days did David spend on this trip? (1 points)

OpenStudy (anonymous):

breaking up the information into parts we have: The company pays $60 a day for food and lodging =$60*x, x is days and $0.65 for each mile traveled. = $0.65*y, y is mile David drove 600 miles, 600 miles = and was reimbursed $3390. $3390 = now we combine what we have learned: y stands for miles and david drove 600 miles, so, 600 = y he was paid $3390 and that money comes from how many miles he drove, $0.65*y, and for how many days he was gone, $60*x. this all becomes $3390 = 0.65*y + 60*x remember y = 600, so $3390 = 0.65*600 + 60*x <-PART A solve for x <-PART B & C

OpenStudy (anonymous):

Thank you for the quick response

OpenStudy (anonymous):

ur welcome. I hope it makes sense